OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 4 times y cubed − the square root of y cubed the square root of y y times the square root of y y times the square root of 3 times y the square root of 3 times y cubed

OpenStudy (dumbcow):

\[\sqrt{4*y^{3}} - \sqrt{y^{3}} = \sqrt{4}\sqrt{y^{3}}-\sqrt{y^{3}} = 2\sqrt{y^{3}}-\sqrt{y^{3}} = \sqrt{y^{3}}\]

OpenStudy (dumbcow):

oh not done yet..haha \[\sqrt{y^{3}} = y^{3/2} = y^{1+1/2} = y*y^{1/2} = y*\sqrt{y}\]